OpenStudy (anonymous):

would someone please answer this>> how to convert from Slope Intercept to Standard Form.

OpenStudy (amistre64):

use algebra

OpenStudy (amistre64):

\[y=mx+c\] \[y=\frac{-a}{b}x+c\] \[by=-ax+c\] \[ax+by=c\]

OpenStudy (amistre64):

and conform c as needed but thats the brunt of it

OpenStudy (amistre64):

lol ... did you ask me if i dont kow the answer?

OpenStudy (anonymous):

So how to get this to standard form? y=3/5x +5

OpenStudy (amistre64):

first step is to get rid of that fraction , mutliply it all by 5

OpenStudy (amistre64):

5(y=3/5x +5) 5y = 3x +25 ; then subtract off the 3x -3x -3x ------------ -3x+5y = 25 ; this is fine but some test dont like a negative out front so change all the signs 3x -5y = -25

OpenStudy (anonymous):

great , thx! you rock!

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

good luck :)